# Valid F-screening in linear regression

### v1

@@ 2.2 A conditional selective inference approach
Step 2 of Box 1 entails a test of $H_{0}^{M}$ if and only if we reject $H_{0}^{1:p}$ (2) at level $\alpha_{0}$: that is, if and only if $Y\in E_{1}$, where
@@
While $E_{1}$ depends on the choice of $\alpha_{0}$, for simplicity in what follows we suppress this.
The left panel of Fig. 4 shows that when restricted to settings where $Y\in E_{1}$ (6), the p-values obtained using $F_{H_{0}^{M}}$ are far from uniform and indeed anti-conservative—i.e., a test based on $p_{H_{0}^{M}}$ (5) does not control the selective Type 1 error in Step 2 of Box 1. This is because $F_{H_{0}^{M}}$ (4) does not follow an $F_{m,n-p-1}$ distribution under $H_{0}^{M}$ when we condition on $Y\in E_{1}$ (6). Our goal is to develop an alternative to $p_{H_{0}^{M}}$ (5) that does lead to selective Type 1 error control.
To illustrate this idea, we consider in Fig. 3 a simple setting with $p=2$ and $\sigma^{2}$ known. The left-hand panel displays the rejection regions for $\chi^{2}$ tests of $H_{0}^{1:2}$ and $H_{0}^{1}$, respectively; this does not account for the sequential nature of the two tests in the context of F-screening. The right-hand panel displays the rejection region of a $\chi^{2}$ test of $H_{0}^{1:2}$, alongside the rejection region of a test of $H_{0}^{1}$ that accounts for the earlier test by conditioning upon rejection of $H_{0}^{1:2}$.
@@
Unfortunately, (7) is not a well-defined probability: it involves nuisance parameters $\left\{\beta_{j}\right\}_{j\notin M}$ not fixed under $H_{0}^{M}$, and thus it cannot be computed. To remedy this, we additionally condition on the event
$E_{2}(y):=\left\{Y:Y^{\top}P_{{X}_{-M}}Y=y^{\top}P_{{X}_{-M}}y\right\};$ (8)
in effect, we hold fixed all the variance in $Y$ that is in the column space of $X_{-M}$. We then define
@@
The next result establishes that a test of $H_{0}^{M}$ based on ${p_{H_{0}^{M}\mid E}}$ controls the selective Type 1 error, conditional on the event that $H_{0}^{1:p}$ was rejected.
@@
A test of $H_{0}^{M}$ based on the p-value ${p_{H_{0}^{M}\mid E}}$ controls the selective Type 1 error, conditional on the event that $H_{0}^{1:p}$ was rejected. That is, for any $\alpha^{\prime}\in(0,1)$,
$\Pr_{\beta_{M}=0}\left({p_{H_{0}^{M}\mid E}}(Y)\leq\alpha^{\prime}\mid Y\in E_{1}\right)=\alpha^{\prime}.$ (11)
The left panel of Fig. 4 empirically corroborates this result. The next result shows that ${p_{H_{0}^{M}\mid E}}$ can be easily approximated via Monte Carlo.
@@
To summarize, a test based on ${p_{H_{0}^{M}\mid E}}$ in (12) controls the selective Type 1 error for $H_{0}^{M}$ in the context of F-screening as described in Box 1 (Theorem 2.1). Furthermore, ${p_{H_{0}^{M}\mid E}}$ can be easily approximated via Monte Carlo, provided that the error variance $\sigma^{2}$ is known (Proposition 2.1). Obtaining confidence intervals for $\beta_{1},\ldots,\beta_{p}$ that attain the nominal selective coverage (conditional on rejecting $H_{0}^{1:p}$), and correcting the corresponding point estimates for selection, are straightforward extensions of the ideas in this section. Details are provided in Appendix B.
@@
In what follows, for conciseness, we may suppress the dependence of some quantities on $y$. E.g., $E(y)$ in (9) may be written $E$, $d(y)$ and $r(y)$ from (13) may be written $d$ and $r$, and ${p_{H_{0}^{M}\mid E}}(y)$ in (12) may be written ${p_{H_{0}^{M}\mid E}}$.
@@ 2.3 A plug-in estimator of the error variance
Proposition 2.1 provides a straightforward way to approximate ${p_{H_{0}^{M}\mid E}}$ (10) via Monte Carlo, provided that $\sigma^{2}$ is known. In practice, we rarely know the true value of $\sigma^{2}$. This situation arises quite often in conditional selective inference, and the typical approach is to use a plug-in estimator for $\sigma^{2}$; see, for example, Lee et al. (2016, §7). Of course, a standard estimator for $\sigma^{2}$ is
@@
Using the notation in Proposition 2.1, this leads us to estimate (10) as
@@
However, we find that ${p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}}$ is quite anti-conservative (see the left panel of Fig. 4).
What has gone wrong? Conditional on the event that we reject $H_{0}^{1:p}$ with ${F_{H_{0}^{1:p}}}$ (3), the estimate $\hat{\sigma}^{2}$ in (14) is biased downward. To understand why, note that there are two reasons we might reject $H_{0}^{1:p}$ using (3): because the numerator is large, or the denominator (which is proportional to $\hat{\sigma}^{2}$) is small. Thus, among datasets for which $H_{0}^{1:p}$ is rejected, $\hat{\sigma}^{2}$ will tend to be small. Furthermore, because (3) and (4) share a denominator, a small value of $\hat{\sigma}^{2}$ will thus be associated with a larger value of (4), leading to a tendency to reject $H_{0}^{M}$ too often. To address this problem, we propose a conditionally “debiased” estimator of the error variance.
@@
Then ${\mathbb{E}}({\tilde{\sigma}^{2}}(\lambda)|Y\in E_{1})=\sigma^{2}$. That is, $\tilde{\sigma}^{2}(\lambda)$ is unbiased for $\sigma^{2}$ conditional on rejection of $H_{0}^{1:p}$. Furthermore, under $H_{0}^{1:p}$, ${\tilde{\sigma}^{2}}:={\tilde{\sigma}^{2}}(0)$ is unbiased for $\sigma^{2}$, conditional on rejection of $H_{0}^{1:p}$.
Under the overall null $H_{0}^{1:p}$ (2), the denominator of ${\tilde{\sigma}^{2}}$ can be easily approximated with Monte Carlo. This leads to the p-value that we use in practice:
@@
We see in the left panel of Fig. 4 that ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$ yields nominal conditional Type 1 error control for values of $\alpha$ that might reasonably be used for hypothesis testing (i.e., $\alpha<0.4).$
From Proposition 2.2, we note that the estimate ${\tilde{\sigma}^{2}}$ used in ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$ is unbiased for $\sigma^{2}$ (conditional on the rejection of $H_{0}^{1:p}$) under $H_{0}^{1:p}$computed under the overall null. Away from the overall null, ${\tilde{\sigma}^{2}}$ is biased upward; however, in our experiments we find that this upward bias is modest, and that its effect on the p-values is minimal (see Appendix C).
3 Computing selective p-values in a retrospective analysis
The left panel of Fig. 4 demonstrates that carrying out Step 2 of Box 1 as though $F_{H_{0}^{M}}$ (4) follows an $F_{m,n-p-1}$ distribution under $H_{0}^{M}$ will lead to failure to control the selective Type 1 error, conditional on having rejected $H_{0}^{1:p}$. However, in the applied literature, examples abound of data analyses where Step 2 was carried out without accounting for Step 1333Here are some examples of published papers that clearly state that the two-step procedure in Box 1 was applied with a standard test in Step 2: (i) “If the groups tested are different, we report the P values of post hoc tests but omit the P value of the omnibus test. If the groups did not test as different, only the P value of the omnibus test is reported” (Kula et al., 2024, page 13); (ii) Table 2 of Wang et al. (2024), where NA’s denote hypotheses that were left untested because the overall test was not rejected; (iii) “Post hoc pairwise comparisons were only performed if the omnibus test was significant” (Tams et al., 2020, caption of Figure 5).. If the original data for such an analysis were available to the reader, then of course they could simply perform a re-analysis with the methods proposed in Section 2. Unfortunately, the raw data are often not available. Therefore, in this section, we show that it is possible to use the usual software output of a least squares linear regression (for instance, the output of summary(lm(y~x)) in R) to “correct” Step 2 of Box 1 in cases where a standard test of $H_{0}^{M}:\beta_{M}=0$ was applied. We refer to this as a retrospective analysis, as it can be applied retrospectively to the results of a published paper, without access to the original data.

### v2

@@ 2.2 A conditional selective inference approach
We test $H_{0}^{M}$ in Step 2 of Box 1 if and only if we reject $H_{0}^{1:p}$ (2) at level $\alpha_{0}\in(0,1)$: that is, if and only if $Y\in E_{1}$, where
@@
The left panel of Fig. 5 shows that conditional on $Y\in E_{1}$ (6), the p-values obtained using $F_{H_{0}^{M}}$ are far from uniform and indeed anti-conservative—i.e., a test based on $p_{H_{0}^{M}}$ (5) does not control the selective Type 1 error in Step 2 of Box 1. This is because $F_{H_{0}^{M}}$ (4) does not follow an $F_{m,n-p-1}$ distribution under $H_{0}^{M}$ when we condition on $Y\in E_{1}$ (6). Our goal is to develop an alternative to $p_{H_{0}^{M}}$ (5) that does lead to selective Type 1 error control.
To illustrate this idea, we consider in Fig. 3 a simple setting with $p=2$ and $\sigma^{2}$ known. The left-hand panel displays the rejection regions for $\chi^{2}$ tests of $H_{0}^{1:2}$ and $H_{0}^{1}$, respectively; this does not account for the sequential nature of the two tests in the context of F-screening. The right-hand panel displays the rejection region of a $\chi^{2}$ test of $H_{0}^{1:2}$, alongside the rejection region of a test of $H_{0}^{1}$ that conditions upon rejection of $H_{0}^{1:2}$.
@@
Unfortunately, (7) is not a well-defined probability: it involves nuisance parameters $\left\{\beta_{j}\right\}_{j\notin M}$ not fixed under $H_{0}^{M}$, and thus it cannot be computed. To remedy this, we additionally condition on the event $E_{2}(y):=\left\{Y:Y^{\top}P_{{X}_{-M}}Y=y^{\top}P_{{X}_{-M}}y\right\};$ in effect, we hold fixed all the variance in $Y$ that is in the column space of $X_{-M}$. We then define
@@
The next result establishes that a test of $H_{0}^{M}$ based on ${p_{H_{0}^{M}\mid E}}(y)$ controls the selective Type 1 error, conditional on the event that $H_{0}^{1:p}$ was rejected.
@@
A test of $H_{0}^{M}$ based on the p-value ${p_{H_{0}^{M}\mid E}}(y)$ (9) controls the selective Type 1 error, conditional on the event that $H_{0}^{1:p}$ was rejected. That is, for any $\alpha^{\prime}\in(0,1)$,
$\Pr_{\beta_{M}=0}\left({p_{H_{0}^{M}\mid E}}(Y)\leq\alpha^{\prime}\mid Y\in E_{1}\right)=\alpha^{\prime}.$
The left panel of Fig. 5 empirically corroborates this result. The next result shows that ${p_{H_{0}^{M}\mid E}}(y)$ can be easily approximated via Monte Carlo.
@@
To summarize, a test based on ${p_{H_{0}^{M}\mid E}}(y)$ in (10) controls the selective Type 1 error for $H_{0}^{M}$ in the context of F-screening as described in Box 1 (Theorem 2.1). Furthermore, ${p_{H_{0}^{M}\mid E}}(y)$ can be easily approximated via Monte Carlo, provided that the error variance $\sigma^{2}$ is known (Proposition 2.1). Obtaining confidence intervals for $\beta_{1},\ldots,\beta_{p}$ that attain the nominal selective coverage (conditional on rejecting $H_{0}^{1:p}$), and correcting the corresponding point estimates for selection, are straightforward extensions of the ideas in this section. Details are provided in Appendix B.
We close this subsection with a result concerning the power of the test that results from using the p-value in (10).
Proposition 2.2.
Consider a fixed sequence of design matrices $X_{n}\in\mathbb{R}^{n\times p}$ and corresponding realizations of the linear model $Y_{n}=X_{n}\beta+\epsilon_{n}$ from (1). Suppose that the design matrices satisfy $\frac{1}{n}X_{n}^{\top}X_{n}\to Q$ for some positive definitive Q. Then the test of $H_{0}^{M}:\beta_{M}=0$ that uses the selective p-value in (10) is consistent under every alternative, conditional on $Y_{n}\in E_{1}$. That is, for any $\beta$ satisfying $\beta_{M}\neq 0$ and arbitrary $\alpha^{\prime}\in(0,1)$, $\lim_{n\to\infty}\;\Pr_{\beta}\left({p_{H_{0}^{M}\mid E}}\left(Y_{n}\right)\leq\alpha^{\prime}|Y_{n}\in E_{1}\right)=1.$
@@
In Proposition 2.2, the following quantities are fixed with respect to $n$: $p\geq 1$, $\beta\in\mathbb{R}^{p}$, $\sigma^{2}>0$, $M\subseteq\{1,\ldots,p\}$. In contrast, $X_{n}$ and $Y_{n}$ grow in length with $n$, and the p-value in (10) depends on $n$ through $Z\sim\chi^{2}_{n-p-1}$, as well as $c$, $d$, and $r$ in (11). In the remainder of the paper, for conciseness, we only include the subscript $n$ when we wish to emphasize dependence on $n$. We may also suppress the dependence of some quantities on $y$ (e.g., ${p_{H_{0}^{M}\mid E}}(y)$ in (10) may be written $p_{H_{0}^{M}\mid E}$}).
2.3 F-Screening and (Conditional) Fisher Information
We now consider the test of $H_{0}^{M}$ in Step 2 of Box 1 through the lens of Fisher information in a simplified setting in which $\sigma^{2}$ is known and $\beta_{0}=0$. Hence, we consider the model (1) without the intercept, and the F-tests in Steps 1 and 2 of Box 1 are replaced with chi-squared tests. The hypothesis of interest in Step 2 is $H_{0}^{1}:\beta_{1}=0$. We compare our proposed approach (specialized to this setting) to a sample splitting procedure.
In the sample splitting procedure, we partition the data into $(X^{\mathrm{tr}},Y^{\mathrm{tr}})$ and $(X^{\mathrm{te}},Y^{\mathrm{te}})$, where a proportion $\rho\in(0,1)$ of the observations are assigned to the training set and the remaining observations are assigned to the test set such that $\rho n$ and $(1-\rho)n$ are integers. We test $H_{0}^{1:p}$ with $(X^{\mathrm{tr}},Y^{\mathrm{tr}})$, and then $(X^{\mathrm{te}},Y^{\mathrm{te}})$ is used to test $H_{0}^{1}$ if and only if $H_{0}^{1:p}$ is rejected.
In the spirit of Fithian et al. (2017, Section 2.5), we derive expressions for: (i) the leftover Fisher information in $Y^{\mathrm{te}}$ about $\beta_{1}$ conditional on the event that $H_{0}^{1:p}$ is rejected with $(X^{\mathrm{tr}},Y^{\mathrm{tr}})$; and (ii) the leftover Fisher information in $Y$ about $\beta_{1}$ conditional on the event that both the overall null $H_{0}^{1:p}$ is rejected using the full data $(X,Y)$ and $P_{X_{-1}}Y=P_{X_{-1}}y$ for some realization $y\in\mathbb{R}^{n}$.
Definition 1 (adapted from Fithian et al., 2017).
Let $\ell(\theta;Y\mid Y\in A)$ denote the conditional log-likelihood of $Y$ given $Y\in A$. The leftover Fisher information about the parameter $\theta$ in data $Y$ conditional on $Y\in A$ is $\mathcal{I}_{Y|Y\in A}(\theta\mid Y\in A):=-{\mathbb{E}}\left[\nabla^{2}\ell(\theta;Y\mid Y\in A)\mid Y\in A\right].$
Proposition 2.3.
  1. i.
    Define $c_{\chi}=\sigma^{2}(\chi^{2}_{p})^{-1}(1-\alpha_{0})$ and the set
    $R_{\rho}(X):=\{Y\in\mathbb{R}^{(\rho n)}:Y^{\top}P_{X}Y\geq c_{\chi}\}.$ (12)
    Treating $\beta_{2},\ldots,\beta_{p}$ as nuisance parameters, the leftover Fisher information in $Y^{\mathrm{te}}$ about $\beta_{1}$ conditional on $Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})$ equals the (unconditional) Fisher information about $\beta_{1}$ in $Y^{\mathrm{te}}$. This quantity is $\frac{1}{\sigma^{2}}\left((X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}\right)$.
  2. ii.
    For any $\tilde{y}\in\mathbb{R}^{n},$ define
    $R(\tilde{y};X):=\{Y\in\mathbb{R}^{n}:Y^{\top}P_{X}Y\geq c_{\chi},\;P_{X_{-1}}Y=P_{X_{-1}}\tilde{y}\}$. (13)
    Treating $\beta_{2},\ldots,\beta_{p}$ as nuisance parameters, the leftover Fisher information in $Y$ about $\beta_{1}$ conditional on $Y\in R(\tilde{y};X)$ equals
    $\mathcal{I}_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};Y\in R(\tilde{y};X))$
    $=\begin{cases}\tfrac{X_{1}^{\top}(I-P_{X_{-1}})X_{1}}{\sigma^{2}}&\text{if }\;\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}\\ \tfrac{X_{1}^{\top}(I-P_{X_{-1}})X_{1}}{\sigma^{2}}\left[1-\left(\tfrac{[\phi(a)-\phi(b)]^{2}}{[1-\Phi(b)+\Phi(a)]^{2}}+\tfrac{\phi(a)\cdot a-\phi(b)\cdot b}{1-\Phi(b)+\Phi(a)}\right)\right]&\text{if }\;\tilde{y}^{\top}P_{X_{-1}}\tilde{y}<c_{\chi}\end{cases}$ (14)
    where
    $a=\tfrac{-\sqrt{c_{\chi}-\tilde{y}^{\top}P_{X_{-1}}\tilde{y}}-\beta_{1}\sqrt{X_{1}^{\top}(I_{n}-P_{X_{-1}})X_{1}}}{\sigma},\quad b=\tfrac{\sqrt{c_{\chi}-\tilde{y}^{\top}P_{X_{-1}}\tilde{y}}-\beta_{1}\sqrt{X_{1}^{\top}(I_{n}-P_{X_{-1}})X_{1}}}{\sigma},$ (15)
    and $\phi$ and $\Phi$ denote the density and distribution functions of the standard normal distribution, respectively.
For insight into the two cases in (14), note that if $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}$, then the set $R(\tilde{y};X)$ in (13) can more succinctly be characterized as $R(\tilde{y};X):=\left\{Y\in\mathbb{R}^{n}:P_{X_{-1}}Y=P_{X_{-1}}\tilde{y}\right\}$, because $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}$ and $P_{X_{-1}}Y=P_{X_{-1}}\tilde{y}$ imply that $Y^{\top}P_{X_{-1}}Y\geq c_{\chi}$ and hence $Y^{\top}P_{X}Y\geq c_{\chi}$. Furthermore, $\tfrac{\partial^{2}}{\partial\beta_{1}^{2}}\ell(\beta_{1};Y\mid P_{X_{-1}}Y=P_{X_{-1}}\tilde{y})=X_{1}^{\top}(I-P_{X_{-1}})X_{1}/\sigma^{2};$ this is simply the (unconditional) Fisher information about $\beta_{1}$ when $\beta_{2},\ldots,\beta_{p}$ are treated as nuisance parameters. Therefore, when $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}$, conditioning on the set $R(\tilde{y};X)$ does not affect the Fisher information about $\beta_{1}$. On the other hand, when $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}<c_{\chi}$, this more succinct characterization of $R(\tilde{y};X)$ is no longer possible, and $\tfrac{\partial^{2}}{\partial\beta_{1}^{2}}\ell(\beta_{1};Y\mid R(\tilde{y};X))\neq X_{1}^{\top}(I-P_{X_{-1}})X_{1}/\sigma^{2}.$ Thus, when $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}<c_{\chi}$, conditioning on $Y\in R(\tilde{y};X)$ indeed affects the Fisher information about $\beta_{1}$.
[…]
$={\mathbb{E}}\left[\mathcal{I}_{Y|Y\in R(\tilde{Y};X)}(\beta_{1};Y\in R(\tilde{Y};X))\mid\tilde{Y}\in R_{1}(X)\right]\cdot\Pr(\tilde{Y}\in R_{1}(X)).$ (16)
The right-hand side of (2.3) is the product of two terms: the expected leftover Fisher information conditional on rejecting $H_{0}^{1:p}$, and the probability of rejecting $H_{0}^{1:p}$.
To approximate the conditional expectation in (2.3), in the left-hand panel of Fig. 4 we generate 1000 draws of $\tilde{Y}$ according to (1) (with $\beta_{0}=0$) and display the average of (14) over all realizations of $\tilde{Y}$ such that $\tilde{Y}^{\top}P_{X}\tilde{Y}\geq c_{\chi}$. The second term on the right-hand side of (2.3) can be computed explicitly by noting that, for $R_{\rho}(X)$ defined in (12) and $\lambda=\frac{1}{2\sigma^{2}}\beta^{\top}X^{\top}X\beta$, $\Pr(Y\in R_{\rho}(X))=1-F_{p,n\cdot\rho-p-1}(\lambda;F^{-1}_{p,n\cdot\rho-p-1}(0;1-\alpha_{0})),$ where $F_{df_{1},df_{2}}(\lambda;\cdot)$ denotes the cumulative distribution function of a noncentral $F$ random variable with $df_{1}$ and $df_{2}$ degrees of freedom and noncentrality parameter $\lambda$. This quantity is displayed in the middle panel of Fig. 4. Finally, in the right-hand panel of Fig. 4 we display the product of the two terms in the left-hand and center panels, i.e., the right-hand side of (2.3).
For the sample splitting procedure, we plot an analogous set of quantities. Recalling from Proposition 2.3(i) that the leftover Fisher information in $Y^{\mathrm{te}}$ about $\beta_{1}$ conditional on $Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})$ is $\frac{1}{\sigma^{2}}\left((X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}\right)$, it follows that
${\mathbb{E}}$ $\left[\tfrac{(X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}}{\sigma^{2}}\cdot\bm{I}_{R_{\rho}(X^{\mathrm{tr}})}(Y^{\mathrm{tr}})\right]=\tfrac{(X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}}{\sigma^{2}}\cdot\Pr(Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})).$ (17)
In the left-hand and center panels of Fig. 4, we display the two terms on the right-hand side of (17), and in the right-hand panel we display their product.
In the left-hand panel of Fig. 4, we see that the selective approach generally results in higher expected leftover Fisher information about $\beta_{1}$, conditional on rejection of $H_{0}^{1:p}$, as compared to sample splitting. Furthermore, the center panel shows that the probability of rejecting $H_{0}^{1:p}$ is always larger using the selective approach than using sample splitting; this is due to the fact that the former uses all $n$ observations whereas the latter uses only $\rho n\leq n$. Finally, the right-hand panel shows that the product of the left and middle panels (representing (2.3) and (17)) is larger for the selective approach than for sample splitting, for all values of $\beta_{1},\ldots,\beta_{p}$ considered.
Remark 2.
We see from (14) that when $\beta_{1}=0$, conditioning on $R(\tilde{Y};X)$ in (13) can increase — and cannot decrease — the leftover Fisher information about $\beta_{1}$ relative to the unconditional Fisher information. This is reflected in the bottom-left panel of Figure 4. However, for larger values of $|\beta_{1}|$, conditioning typically does not increase the leftover Fisher information.
$\beta_{1}=0.5$ $\beta_{1}=0.25$ $\beta_{1}=0$
   $\mathbb{E}$$\big$[Leftover Fisher $\mid\text{rej }\bm{H_{0}^{1:p}}\big]$      Pr(rej $\bm{H_{0}^{1:p}}$)       $\mathbb{E}$$\big$[Leftover Fisher $\mid\text{rej }\bm{H_{0}^{1:p}}\big]\bm{\cdot}$ Pr(rej $\bm{H_{0}^{1:p}}$)
Figure 4: We construct an $n\times p$ orthogonal design matrix $X$ with $n=100$ and $p=5$. Then, for $\beta_{1}\in\{0.5,0.25,0\}$ and for a range of values of $s\in[-1,1]$, we generate 1000 response vectors $\tilde{Y}$ according to (1) with $\sigma=1$, $\beta_{j}=s$ for all $j\in\{2,\ldots,p\}$, and $\beta_{0}=0$. We then compute quantities related to the leftover Fisher information about $\beta_{1}$ for our selective approach, and for sample splitting with split proportions $\rho\in\{0.1,0.5,0.9\}$. Left: For each of the 1000 realizations $\tilde{y}$ of $\tilde{Y}$, we compute the leftover Fisher information $\mathcal{I}_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};Y\in R(\tilde{y};X))$ for our conditional selective inference procedure and display the average over all datasets such that $\tilde{Y}\in R_{1}(X)$. This approximates ${\mathbb{E}}\left[\mathcal{I}_{Y\mid Y\in R(\tilde{Y};X)}(\beta_{1};Y\in R(\tilde{Y};X))\mid\tilde{Y}\in R_{1}(X)\right]$. Further, for $\rho\in\{0.1,0.5,0.9\}$, we display $(X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}/\sigma^{2}$. Center: We display $\Pr(Y\in R_{1}(X))$, along with $\Pr(Y^{\mathrm{tr}}_{\rho}\in R_{\rho}(X^{\mathrm{tr}}))$ for $\rho\in\{0.1,0.5,0.9\}$. Right: We display the product of the left-hand and center columns, i.e., the right-hand sides of (2.3) and (17).
@@ 2.4 A plug-in estimator of the error variance
Proposition 2.1 provides a straightforward way to approximate $p_{H_{0}^{M}\mid E}$} (9) via Monte Carlo, provided that $\sigma^{2}$ is known. In practice, we rarely know the true value of $\sigma^{2}$. The typical solution is to use a plug-in estimator for $\sigma^{2}$; see, for example, Lee et al. (2016, §7). A standard estimator for $\sigma^{2}$ is
@@
Using the notation in Proposition 2.1, this suggests we might estimate (9) as
@@
However, $p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}$} is quite anti-conservative (see the left panel of Fig. 5).
What has gone wrong? Conditional on the event that we reject $H_{0}^{1:p}$ with $F_{H_{0}^{1:p}}$} (3), the estimate $\hat{\sigma}^{2}$ in (18) is biased downward. To understand why, note that there are two reasons we might reject $H_{0}^{1:p}$ using (3): because the numerator is large, or the denominator (which is proportional to $\hat{\sigma}^{2}$) is small. Thus, among datasets for which $H_{0}^{1:p}$ is rejected, $\hat{\sigma}^{2}$ will tend to be small. Furthermore, because (3) and (4) share a denominator, a small value of $\hat{\sigma}^{2}$ is associated with a larger value of (4), leading to a tendency to reject $H_{0}^{M}$ too often. To address this problem, we propose a conditionally “debiased” estimator of the error variance.
@@
Then ${\mathbb{E}}({\tilde{\sigma}^{2}}(\lambda)|Y\in E_{1})=\sigma^{2}$, i.e., ${\tilde{\sigma}^{2}}(\lambda)$ is unbiased for $\sigma^{2}$ conditional on rejection of $H_{0}^{1:p}$.
An immediate consequence of Proposition 2.4 is that under $H_{0}^{1:p}$ (2), ${\tilde{\sigma}^{2}}:={\tilde{\sigma}^{2}}(0)$ is unbiased for $\sigma^{2}$ conditional on rejection of $H_{0}^{1:p}$. Further, under $H_{0}^{1:p}$, the denominator of $\tilde{\sigma}^{2}$} can be easily approximated with Monte Carlo. This leads to the p-value that we use in practice:
@@
Away from the overall null, $\tilde{\sigma}^{2}$} is biased upward; however, in our experiments we find that this upward bias is modest, and that its effect on the p-values is minimal (see Appendix C). We see in the left panel of Fig. 5 that $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} yields nominal selective Type 1 error control for values of $\alpha$ that might reasonably be used for hypothesis testing (i.e., $\alpha<0.4).$ In the next proposition we establish that $\tilde{\sigma}^{2}$} is consistent for $\sigma^{2}$, conditional on $Y\in E_{1}$.
Proposition 2.5.
Consider the same setting as described in Proposition 2.2 and Remark 1, but without any requirement on $\beta$. Then conditional on $Y_{n}\in E_{1},$ the estimator ${\tilde{\sigma}^{2}}={\tilde{\sigma}^{2}}(0)$ (20) converges in probability to $\sigma^{2},$ i.e., $\forall\varepsilon>0$, $\lim_{n\to\infty}\Pr\left(\left|{\tilde{\sigma}^{2}}-\sigma^{2}\right|>\varepsilon\mid Y_{n}\in E_{1}\right)=0$.
Consistency of $\tilde{\sigma}^{2}$} for $\sigma^{2}$ conditional on $Y_{n}\in E_{1}$ allows us to establish that ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}(Y_{n})$ controls the selective Type 1 error asymptotically.
Proposition 2.6.
Under $H_{0}^{M}:\beta_{M}=0$ and conditional on $Y_{n}\in E_{1}$, the p-value $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} in (21) converges to the $\text{Unif}(0,1)$ distribution as $n\to\infty$. That is, for all $\alpha^{\prime}\in(0,1)$, $\lim_{n\to\infty}\Pr_{\beta_{M}=0}\left({p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}(Y_{n})\leq\alpha^{\prime}\mid Y_{n}\in E_{1}\right)=\alpha^{\prime}$.
3 Selective p-values in a retrospective analysis
In the applied literature, examples abound of analyses where Step 2 was carried out without accounting for Step 1333Here are some examples of published papers that clearly state that the two-step procedure in Box 1 was applied with a standard test in Step 2: (i) “If the groups tested are different, we report the p-values of post hoc tests but omit the p-value of the omnibus test. If the groups did not test as different, only the P value of the omnibus test is reported” (Kula et al., 2024, page 13); (ii) Table 2 of Wang et al. (2024), where NA’s denote hypotheses that were left untested because the overall test was not rejected; (iii) “Post hoc pairwise comparisons were only performed if the omnibus test was significant” (Tams et al., 2020, caption of Figure 5).. If the original data for such an analysis were available to the reader, then they could reanalyze the data using the methods in Section 2. Unfortunately, the raw data are often not available. In this section, we demostrate that one can use the usual software output of a least squares linear regression (e.g., the output of summary(lm(y~x)) in R) to “correct” Step 2 of Box 1 when a standard test of $H_{0}^{M}$ was used. We refer to this as a retrospective analysis, as it can be applied retrospectively to the results of a published paper, without access to the raw data.